OpenStudy (anonymous):

The average of 6 numbers is 13. Two of the numbers are 10 and 16. What is the average of the remaining four numbers

OpenStudy (anonymous):

http://www.algebra.com/algebra/homework/Average/Average.faq.question.323608.html very same exercise

OpenStudy (anonymous):

Different values tho

OpenStudy (anonymous):

10+16+\[\frac{10+16+x+y+z+w}{6}=13\]

OpenStudy (anonymous):

\[x+y+z+w=52\]

OpenStudy (anonymous):

\[x+y+z+w/4 \] is the average of four numbers..... which is 13

OpenStudy (anonymous):

Calculate the total sum of the six numbers, substracting the numbers you know you find the sum of the other four numbers. The rest you do it (how to calculate the average of four numbers?)